OpenStudy (anonymous):

Consider a line tangent to the curve x^(2/3)+y^(2/3)=1. Show that the distance between the x-intercept and the y-intercept of this line is always 1.

OpenStudy (anonymous):

Are you sure about your equation?

OpenStudy (anonymous):

Your equation is fine. Stay tuned for the answer.

OpenStudy (anonymous):

yes i'm sure.it is \[x^{2/3}+y \left\{ 2/3\right\}\]

OpenStudy (anonymous):

Is it \[ x^{2/3} + y^{2/3}=1 \]

OpenStudy (anonymous):

I have trig solution for it

OpenStudy (anonymous):

yes it is. what is your solution ?

OpenStudy (anonymous):

Put \[ x(t)=\cos ^3(t)\\ y(t)=\sin^3(t) \]

OpenStudy (anonymous):

Let c be a value of t where we draw the tangent. The slop at c is \[ \frac{y'(c)}{x'(c)}=-\tan(c) \]

OpenStudy (anonymous):

You do the computation

OpenStudy (anonymous):

The equation of the tangent at t=c is \[ y- y(c) = -\tan(c)(x-x(c)) \]

OpenStudy (anonymous):

Put x=0 and find the y-intercept b to be \[ b=\sin ^3(c)+\sin (c) \cos ^2(c) \]

OpenStudy (anonymous):

Put y=0 and find the x-intercept a to be \[ a=\cot (c) \left(\sin ^3(c)+\sin (c) \cos ^2(c)\right) \]

OpenStudy (anonymous):

Compute \[ a^2 + b^2 \] and you find that is equal to 1

OpenStudy (anonymous):

What is your background in Math?

OpenStudy (anonymous):

i have a question. how do you find the slope that is -tan(c)

OpenStudy (anonymous):

I'm first year student in univercity

OpenStudy (anonymous):

I will send you a solution that does not involve trig

OpenStudy (anonymous):

Is this part of your implicit differentiation subject?

OpenStudy (anonymous):

yes it is. i couldn't understand the subject well